Please help.
I am in the dark and I am counting my change. I can't turn on the light so I try to figure out how much money I have. I know that $2.50 in change, but I am going to the laundry so I need quarters. I remember that I have 3 dimes and that the rest are quarters and nickels. I count my money, and I have 19 coins. How many quarters and how many nickels do I have?
Thank you for your time. ☺

Answer by Nate(3500)

we will say that $2.50 is 250
we will say that $0.10 is 10
we will say that $0.25 is 25
250 is change minus 30 in dimes
=220
define variables:
n=nickels and q=quarters
n+q=16
(5)n+(25)q=220
if you combine, you get 7 coins for quarters and 9 coins in nickels and 3 coins for dimes